OpenStudy (anonymous):

What is the greatest common factor of 52x5 and 12x4? A. 4x4 B. 4x5 C. 6x4 D. 6x5

OpenStudy (calculusxy):

First compute of what is the GCF of 52 and 12.

OpenStudy (anonymous):

The GCF for 52 and 12 is 4 and the GCF for 5 and 4 is 1

OpenStudy (anonymous):

do the prime factorization of each an it shoud be easy to see. it doesn't look as if your choices contain the actual answer, though.

OpenStudy (anonymous):

do the prime factorization of 5 and 1?

OpenStudy (calculusxy):

If I'm not wrong, do the prime factorisation of 5 and 4.

OpenStudy (anonymous):

52 x 5 = 2*2*13*5 12 x 4 = 2*2*2*2*3 What's the gcd?

OpenStudy (anonymous):

You guy's are confusing me !

OpenStudy (calculusxy):

@pgpilot326 Can you explain the equation after 52x5 and so forth

OpenStudy (anonymous):

@calculusxy ?

OpenStudy (anonymous):

Do you mean: What is the greatest common factor of \(52x^5\) and \(12x^4\)? A. \(4x^4\) B. \(4x^5\) C. \(6x^4\) D. \(6x^5\)

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

Yeah that's how its written but i copy and paste on here so it comes out like that

OpenStudy (calculusxy):

Then it should be 4x4

OpenStudy (anonymous):

Next time, use the equation editor to avoid confusion. The button is located on the bottom of the reply box.

OpenStudy (anonymous):

k

OpenStudy (anonymous):

that'd change things a little... \[52x ^{4}= 2*2*13*x*x*x*x*x,12x ^{4}=2*2*3*x*x*x*x\]